### v1

@@ 1 Introduction
Finding approximate fixed points is a well-studied problem in the case of the $\ell_{2}$-metric and the $\ell_{\infty}$-metric (see e.g. [sikorskiComputationalComplexityFixed2009, chenComputingFixedPoint2024]). In particular, efficient algorithms in the case of the $\ell_{2}$-metric have been known since 1993 [sikorskiEllipsoidAlgorithmComputation1993], and only recently the first polynomial query upper bound for the $\ell_{\infty}$-metric was found by Chen et al. [chenComputingFixedPoint2024]. They also mention the case of the $\ell_{1}$-metric as an open problem, albeit without mentioning any applications. To the best of our knowledge, our reduction from G-Arrival to $\ell_{1}$-contraction provides the first concrete application for this problem, further motivating the study of its complexity. We think that our reduction could eventually lead to better algorithms for Arrival through the study of $\ell_{1}$-contraction maps.
@@ 1.1 Related Work
It is also known that Arrival can be solved in polynomial time on some restricted graph classes. For example, a result due to Priezzhev et al. [priezzhevEulerianWalkersModel1996] implies that Arrival can be solved by simulation in polynomial time on Eulerian graphs. Other results include polynomial-time algorithms on tree-like multigraphs [augerPolynomialTimeAlgorithm2022] and path-like multigraphs with many tokens [augerGeneralizedARRIVALProblem2023].
@@ 2.1 Treewidth and Balanced Separators
If $G$ has treewidth at most $k$, then for any $S\subseteq V$, the subgraph $G-S$ has a balanced separator of size at most $k+1$.
In other words, this lemma allows us to use the treewidth of $G$ to infer the existence of small balanced separators in all induced subgraphs of $G$.

### v2

@@ 1 Introduction
Finding approximate fixed points is a well-studied problem in the case of the $\ell_{2}$-metric and the $\ell_{\infty}$-metric (see e.g. [sikorskiComputationalComplexityFixed2009, chenComputingFixedPoint2024]). In particular, efficient algorithms in the case of the $\ell_{2}$-metric have been known since 1993 [sikorskiEllipsoidAlgorithmComputation1993], and Chen et al. [chenComputingFixedPoint2024] only recently gave the first polynomial query upper bound for the $\ell_{\infty}$-metric. Even more recently, a generalization of the result by Chen et al. to $\ell_{p}$-contractions for every $p\in[1,\infty]$ was announced [haslebacherQueryEfficientFixpoints\$ell_p\$Contractions2025]. Concretely, this means that an approximate fixed point of an $\ell_{1}$-contraction can be found with polynomially many queries to the contraction map.
Unfortunately, the algorithm for $\ell_{1}$-contractions in [haslebacherQueryEfficientFixpoints\$ell_p\$Contractions2025] is only query-efficient (and not time-efficient). Thus, our reduction currently does not imply better algorithms for G-ARRIVAL. We still think that our reduction is interesting: To the best of our knowledge, it provides the first concrete application for the $\ell_{1}$-contraction problem. Moreover, the polynomial query upper bound for $\ell_{1}$-contractions [haslebacherQueryEfficientFixpoints\$ell_p\$Contractions2025] suggests that better (maybe even time-efficient) algorithms could be obtained for $\ell_{1}$-contraction and hence G-ARRIVAL in the future. For this, it might also be interesting to observe that the $\ell_{1}$-contraction map obtained through our reduction from G-ARRIVAL has the additional property of being monotone with respect to the coordinate-wise partial order. Given recent work of Batziou et al. [batziouMonotoneContractions2024] on monotone $\ell_{\infty}$-contractions, it seems plausible that $\ell_{1}$-contraction and monotonicity could maybe be exploited simultaneously as well.
@@ 1.1 Related Work
It is also known that ARRIVAL can be solved in polynomial time on some restricted graph classes. For example, a result due to Priezzhev et al. [priezzhevEulerianWalkersModel1996] implies that ARRIVAL can be solved by simulation in polynomial time on Eulerian graphs. Other results include polynomial-time algorithms on tree-like multigraphs [augerPolynomialTimeAlgorithm2022] and path-like multigraphs with many tokens [augerGeneralizedARRIVALProblem2023]. More recently, algorithms running in quasi-polynomial-time were given for the case of tree-like multigraphs with many tokens [ghorbaniQuasiPolynomialTimeAlgorithm2025]. Our algorithm effectively generalizes this result to all graphs of bounded treewidth.
@@ 2.1 Treewidth and Balanced Separators
If $G$ has treewidth at most $k$, then for every $S\subseteq V$, the subgraph $G-S$ has a balanced separator of size at most $k+1$.
This lemma allows us to use the treewidth of $G$ to infer the existence of small balanced separators in all induced subgraphs of $G$.